An open economy with a government sector is in equilibrium. Assume the following:  Marginal propensity to save = 0.4

Marginal propensity to tax = 0.2

Marginal propensity to import = 0.2

Showing your method of working, calculate by how much the equilibrium level of national income would fall, if injections in the economy are reduced by $60m.


1
Expert's answer
2021-11-08T11:41:26-0500

N.I(Y)=C+I-M


I)We first calculate after tax income:


Marginal propensity to tax = 0.2


Therefore,tax amount is 0.2Y and after tax income is Y-0.2Y=0.8Y


II)We then calculate consumption:


Marginal propensity to save = 0.4 therefore the marginal propensity to consume becomes 1-0.4=0.6


Therefore consumption C=0.6"\times" 0.8Y=0.48Y


III)We then calculate Imports;

=after tax income "\times" Marginal propensity to import


=0.8Y"\times" 02


= 0.16Y


N.I(Y)=C+I-M

Inputting the values obtained above:


Y=0.48Y+"\$60"-0.16Y

0.68Y="\$60"

Y="\$" 88.24


Therefore if government injections reduced by "\$60m" ,N.I equilibrium would reduce by:


"\frac{\$88.25-\$60}{\$88.25}\times 100=32\%"